OpenStudy (anonymous):

r times the square root of r − 3 times r times the square root of r + 4 times r times the square root of r

OpenStudy (anonymous):

\[r*\sqrt{r-3}*r*\sqrt{r+4}*r*\sqrt{r}\]\[r^3r^{1/2}(r-3)^{1/2}(r+4)^{1/2}\]\[r^{7/2}(r-3)^{1/2}(r-4)^{1/2}\]. That's as simplified as it gets. The roots are 0, 3, and 4.

OpenStudy (anonymous):

@mitodoteira You beat me to answering

OpenStudy (anonymous):

Thnx!